OpenStudy (anonymous):

Is 216 divisble by 3 A.no B.yes,because the sum 2+1+6=9 which is divisible by 3 C. yes,because the last digit of the number is 0 Plz hurry

OpenStudy (anonymous):

B.

OpenStudy (anonymous):

yes because the sum 2+1+6=9 which is divisible by 3

OpenStudy (anonymous):

okay i have another question to ask

OpenStudy (anonymous):

Is 26 divisible by 2 No Yes because the sum of 2+1+6+9,which is divisible by 3 Yes because the last digit is 6 which is divisible by 3

OpenStudy (anonymous):

yes it is divisible by 2 but im not sure why

OpenStudy (anonymous):

i thought because it is an even number

OpenStudy (anonymous):

okay which answer choice A,B,or C

OpenStudy (anonymous):

it is not divisible by 3 because a number must equal up to a number divisible by 3 to be divisible by 3. for example 369 is divisible by 3 because 3+6+9=18. 1+8=9 and 9/3=3

OpenStudy (anonymous):

i dont know sorry

OpenStudy (anonymous):

OKkay can u healp me with 9 question

OpenStudy (anonymous):

9 question?

OpenStudy (anonymous):

i would say C but im not sure